# Universal quantum control over Majorana zero modes

### v1

@@ Abstract
Majorana zero mode (MZM) exhibits inherent resilience to local parametric fluctuations in quantum information processing, due to a topological protection mechanism in the non-Abelian braiding statistics of the anyonic quasiparticles. In this paper, we construct the braiding operations between arbitrary pair of three MZMs under the framework of universal quantum control (UQC). Largely detuned driving fields on the mediator, a local defect of lattice, enable indirect and tunable exchange interaction among arbitrary two MZMs. The braiding operations can then present through the evolution along the universal nonadiabatic passages, whose robustness against the driving-field errors can be substantially enhanced by rapid modulation of the passage-dependent global phase. Moreover, a chiral population transfer on MZMs along the universal passage can be perfectly demonstrated in both clockwise and counterclockwise manners. Our scale-free protocol therefore provides an avenue towards the universal quantum control over MZMs, which is fundamental and essential for topological quantum computation.
@@ I Introduction
Certain problems Bravyi et al. (2018) can show provable advantages of quantum computer over its classical counterpart, i.e., Shor’s algorithm for factoring large numbers Calderbank and Shor (1996); Ekert and Jozsa (1996); Kivlichan et al. (2020) and Grover’s algorithm for amplifying probability of target states Grover (1997); Nagib et al. (2025). However, practical quantum computer demands large-scale quantum systems with extensive gate operations Gidney and Ekerå (2019). Each elementary operation can introduce systematic errors that accumulate exponentially, posing significance of error correction schemes, such as the Shor’s code Shor (1995). That scheme encodes a single logical qubit into nine physical qubits and can detect and correct single-qubit errors. Another paradigm for error correction is topological quantum computation Nayak et al. (2008); Freedman (1998); Kitaev (2001, 2003); Das Sarma et al. (2005); Stern and Halperin (2006); Bonderson et al. (2006), in which qubits are encoded on the non-Abelian anyonic quasiparticles Nayak et al. (2008). Quantum information encoded in those non-local degrees of freedom is believed to maintain intrinsic fault-tolerance against local parametric fluctuations.
@@
Anyons fundamentally bifurcate into the Abelian Wilczek (1990) and non-Abelian categories Read and Green (2000); Ivanov (2001); Alicea et al. (2011); Sarma et al. (2015). The unitary evolutions resulted from braiding Abelian anyons yield the accumulation of phases: the wave function of multiple Abelian anyons acquires a fractional phase factor $e^{i\theta}$, $\theta\neq 0,\pi$, on the exchange of quasiparticles. Bosons ($\theta=0$) and fermions ($\theta=\pi$) are the limiting cases Wilczek (1990). Besides the phase accumulation, the braid group Ivanov (2001); Alicea et al. (2011); Sarma et al. (2015) associated with the non-Abelian anyons enables nontrivial unitary operations among the degenerate many-quasiparticle states. Due to the fact that the qubits encoded in the global degenerate states are robust against the local systematic errors Witzel and Das Sarma (2006), the non-Abelian anyons can serve as the primitive units of topological quantum computation.
@@
Majorana zero mode (MZM), first theoretically predicted in the Kitaev chain model Kitaev (2001), represents the most elementary realization of the non-Abelian anyons. The Kitaev chain models a one-dimensional system of spinless fermions and was once regarded as a somewhat unphysical toy model Elliott and Franz (2015). Nevertheless, the spinless fermions behavior can be simulated through experimentally accessible proxy systems, such as the spin-orbit-coupling electron gases under Zeeman field, as suggested by the subsequent theoretical studies Elliott and Franz (2015); Fu and Kane (2008); Sau et al. (2010); Alicea (2010); Lutchyn et al. (2010); Oreg et al. (2010); Nadj-Perge et al. (2013); Pientka et al. (2013); Gao et al. (2024). Experimental researches for MZM have been conducted in multiple physical platforms, including the superconducting systems Nadj-Perge et al. (2014); Kim et al. (2018); Jiao et al. (2018); Schneider et al. (2020); Yang et al. (2021) and hybrid magnetic-superconductor systems Schneider et al. (2022). Recently, indirect evidences for MZMs have been witnessed in the hybrid semiconductor-superconductor system Quantum (2025), exhibiting a zero-bias conductance peak Das et al. (2012); Deng et al. (2012); Finck et al. (2013); Churchill et al. (2013) and the topological protection against the variations of local parameters Jiao et al. (2018); Yang et al. (2021); Quantum (2025). The braiding of MZMs, however, remains the gold-standard signature for definitively establishing their non-Abelian statistics Dvir et al. (2023).
@@
Conventional braiding operations of MZMs were proposed under the adiabatic condition Ivanov (2001); Stone and Chung (2006); Read (2009); Sau et al. (2011); van Heck et al. (2012); Kraus et al. (2013) and governed by the Berry-Wilczek-Zee holonomy Wilczek and Zee (1984); Cheng et al. (2011). In practice, the implementation of topological quantum computation based on the adiabatic braiding operations is constrained by the extended operation time. First, the prolonged exposure to the environment can induce ruinable decoherence effect on the quantum systems Carmichael (1999). Second, while the exact adiabatic solution is established within infinite evolution time, the realistic operation with a finite clock speed unavoidably entails nonzero non-adiabatic errors Nayak et al. (2008). Such errors will lead to the leakage from the subspace of degenerate ground states to the external Hilbert space. Third, the long operation time of topological quantum computer will increase the possibility of the quasiparticle poisoning Catelani et al. (2011); Karzig et al. (2021); Liu et al. (2024), i.e., a quasiparticle enters the collection of MZMs and changes the parity of the states. Although the measurement-based tools Knapp et al. (2016) have been used to mitigate unwanted transition errors during adiabatic braid operations Bonderson et al. (2008); Bomantara and Gong (2020), a fast and high-fidelity nonadiabatic braiding operation remains desired for MZMs.
@@
Inspired by transitionless quantum driving Berry (2009), the shortcuts to non-Abelian braiding Karzig et al. (2015); Song et al. (2025) exploit the time-dependent and direct exchange interaction between MZMs as the counterdiabatic driving fields. But practical implementation of such interactions in semiconductor-superconductor systems is fundamentally difficult Amundsen et al. (2024). In analog to nonadiabatic holonomic quantum computation Sjöqvist et al. (2012); Liu et al. (2019), the nonadiabatic braiding operations Yu et al. (2025) have been demonstrated between two MZMs mediated by the lattice defect (LD). However, the braiding operations under the parallel-transport condition Jing et al. (2017) is sensitive to the systematic errors of the local driving fields on LD.
@@
We here construct the braiding operations between any pair of three MZMs mediated with LD under local driving fields by universal quantum control (UQC) Jin and Jing (2025a, b, c, d). It is a theoretical framework that unifies the shortcuts to adiabaticity and holonomic quantum transformation and is beyond the parallel-transport condition. The degree of freedom of LD can be eliminated Jin and Jing (2025d); James and Jerke (2007) when the driving fields are largely detuned; and then it enables indirect exchange interaction among MZMs. Consequently, the braiding operations of arbitrary two MZMs can be performed along the universal passages of the system. Our protocol presents fault-tolerant braiding operations that are robust against significant fluctuations in both transition frequency of LD and Rabi frequency of driving field. In addition, the clockwise and counterclockwise population transfers of three MZMs can be demonstrated by the time evolution along the passage that is superposed by LD excited state and MZMs.
@@
The rest part of this paper is structured as follows. In Sec. II.1, we introduce a general model that is consisted of three MZMs coupled to a lattice defect and the full Hamiltonian. In Sec. II.2, we derive the effective Hamiltonian by using the local largely-detuned driving fields on LD, based on which the braiding operations can be established via universal nonadiabatic passages. In Sec II.3, we employ a rapidly-varying global phase to correct systematic errors. Fault-tolerant, fast, and high-fidelity braiding operations can then be constructed for MZMs. Section III devotes to designing the chiral population transfer of three MZMs in both clockwise and counterclockwise manners. The whole work is concluded in Sec. IV. Appendix A presents the UQC theory in the parametric space with ideal control setting, followed by an error-correction extension in Appendix B.
@@ II.1 Model and Hamiltonian
Figure 1: (a) Sketch of an atomic chain system comprising a LD and three one-dimensional NN-site Kitaev chains. LD is coupled to MZMs $\gamma_{1,1}$, $\gamma_{1,2}$, and $\gamma_{1,3}$ through local off-resonant driving fields applied to the LD. Inset: the four-level transition diagram for the effective subspace of LD and MZMs. (b) Effective two-level transition diagram for the braiding operation between MZMs $\gamma_{1,j}$ and $\gamma_{1,k}$, $j\neq k$.
Our braiding protocol is conducted on the atomic-chain system as shown in Fig. 1(a). The system comprises a two-level system (lattice defect) as a mediator, consisting of an excited state $|e\rangle$ and a ground state $|g\rangle$, and three one-dimensional Kitaev chains Kitaev (2001, 2003) of open boundary condition, each of which is composed of $N$ sites. The system Hamiltonian can be written as Kitaev (2001, 2003) ($\hbar\equiv 1$)
$H(t)$ $=\sum_{k=1}^{3}\Big{[}-\mu\sum_{n=1}^{N}c_{n,k}^{\dagger}c_{n,k}-\sum_{n=1}^{N-1}(Jc_{n+1,k}^{\dagger}c_{n,k}+$ (1)
@@
where $\sigma_{+}\equiv|e\rangle\langle g|$ ($\sigma_{-}=|g\rangle\langle e|$) and $c_{n,k}^{\dagger}$ ($c_{n,k}$) are the creation (annihilation) operator for LD and the $n$th site of the $k$th Kitaev chain, respectively. The fermions in the Kitaev chains are assumed to have the same transition frequency $\mu$. $\omega(t)$ is the transition frequency of the two-level lattice defect, which is set as a constant in our protocol despite it can be tuned via an external magnetic field Wolfowicz et al. (2021). $J$ and $g$ are the hopping and pairing amplitudes between the fermions in the chains, respectively.
The Hamiltonian in Eq. (1) admits an alternative representation in terms of Majorana fermion quasiparticles, which emerge as bounded pairs in the system. Specifically, the Majorana fermions of the $k$th Kitaev chain, $1\leq k\leq 3$, can be defined as Kitaev (2001, 2003)
@@
where $n$ ranges from $1$ to $N$. As defined in Eq. (2), Majorana fermions are characterized by two fundamental properties: (i) self-conjugation, $\gamma_{n,k}^{\dagger}=\gamma_{n,k}$ and (ii) anti-commutation relations $\{\gamma_{n,j},\gamma_{m,k}\}=2\delta_{nm}\delta_{jk}$, where $n$ and $m$ range from $1$ to $2N$. Under the parametric conditions of $\mu=0$ and J=gJ=-g Kitaev (2001, 2003) and Eq. (2), the Hamiltonian (1) is transformed to be
@@
Equations (2) and (3) indicate that there exist two decoupled Majorana zero modes in each Kitaev chain, i.e., $\gamma_{1,k}$ and $\gamma_{2N,k}$.
We are desired to construct the braiding operations among chosen MZMs $\gamma_{1,j}$ and $\gamma_{1,k}$, $j\neq k$, via a mediator LD. In the absence of LD, when the system is initially in the vacuum state $|0\rangle$, the dynamics is constrained in the subspace spanned by the states $|\gamma_{1,k}\rangle=\gamma_{1,k}|0\rangle$. In the presence of the coupled LD under local driving fields, the full Hamiltonian can be expressed as Yu et al. (2025)
@@
with the driving Hamiltonian
$H_{d}(t)=2\sum_{k=1}^{3}\Omega_{k}(t)\cos\left[\int_{0}^{t}\omega_{k}(s)ds+\varphi_{k}\right]|e\rangle\langle\gamma_{1,k}|+{\rm H.c.},$ (5)
where $\Omega_{k}(t)$, $\omega_{k}(t)$, and $\varphi_{k}$ are the Rabi frequency, the driving frequency, and the phase of the driving fields, respectively. While $\gamma_{1,k}$’s involve only a few sites of the Kitaev chains, the driving Hamiltonian in Eq. (5) can be engineered through local pairing and hopping couplings between LD and the proximal sites of the chains. It has been extensively studied, both theoretically Liu et al. (2013); Li et al. (2014); Long et al. (2021); Nathan et al. (2021); Xu et al. (2023) and experimentally Cooper et al. (2019); Miles et al. (2025), in cold atoms and quantum-dot arrays.
Adapt from MZMs, the full Hamiltonian $H_{\rm rot}(t)$ (4) indicates that there are no interactions between LD and the rest Majorana fermions $\gamma_{2n}$ and $\gamma_{2n+1}$, $1\leq n\leq N-1$. Then we can focus on
@@
$\times\cos\left[\int_{0}^{t}\omega_{k}(t^{\prime})dt^{\prime}+\varphi_{k}\right]|e\rangle\langle\gamma_{1,k}|+{\rm H.c.}\Big{\}}.$
@@
where the detuning $\Delta_{k}(t)\equiv\omega_{k}(t)-\omega$. The relevant level transition diagram can be found in the inset of Fig. 1(a).

### v2

@@ Abstract
Majorana zero mode (MZM) exhibits inherent resilience to local parametric fluctuations, due to the topological protection mechanism in the non-Abelian braiding statistics of the anyonic quasiparticles. In this paper, we construct the braiding operations between an arbitrary pair in three MZMs under the theoretical framework of universal quantum control. Largely detuned driving fields on the mediator, a local defect of lattice, enable indirect and tunable exchange interaction between arbitrary two MZMs. The braiding operations can then emerge through the time evolution along the universal nonadiabatic passages, whose robustness against driving-field errors and quasiparticle poisoning can be substantially enhanced by the rapid modulation over the passage-dependent global phase. Moreover, a chiral transfer for population on MZMs along the universal passage can be perfectly demonstrated in both clockwise and counterclockwise manners without eliminating the mediator defect. Our protocol is linear scalable and provides an avenue towards the universal quantum control over MZMs, which is fundamental and essential for topological quantum computation.
@@ I Introduction
Certain problems Sergey2018Quantum can show provable advantages of quantum computer over its classical counterpart, i.e., Shor’s algorithm for factoring large numbers Calderbank1996Good; Ekert1996Quantum; Kivlichan2020Improved and Grover’s algorithm for amplifying probability of target states Grover1997; Nagib2025. However, practical quantum computer demands large-scale quantum systems with extensive gate operations Gidney2019HowTF. Each elementary operation can introduce systematic errors that accumulate exponentially, posing significance of error correction schemes, such as the Shor’s code Shor1995Scheme. That scheme encodes a single logical qubit into nine physical qubits and can detect and correct single-qubit errors. Another paradigm for error correction is topological quantum computation Nayak2008NonAbelian; Michael1998PNP; Kitaev2001Unpaired; Kitaev2003Fault; Das2005Topologically; Stern2006Proposed; Bonderson2006Detecting, in which qubits are encoded in the non-Abelian anyonic quasiparticles Nayak2008NonAbelian. Quantum information encoded in those nonlocal degrees of freedom is believed to maintain intrinsic fault tolerance against local parametric fluctuations.
@@
Anyons fundamentally bifurcate into the Abelian Wilczek1999Fractional and non-Abelian categories Read2000Paired; Ivanov2001NonAbelian; Alicea2011NonAbelian; Sarma2015Majorana. The unitary evolutions resulting from braiding Abelian anyons yield the accumulation of phases: the wave function of multiple Abelian anyons acquires a fractional phase factor $e^{i\theta}$ with $\theta\neq 0,\pi$ on the exchange of quasiparticles. Bosons ($\theta=0$) and fermions ($\theta=\pi$) are their limiting cases Wilczek1999Fractional. Besides the phase accumulation, the braid group Ivanov2001NonAbelian; Alicea2011NonAbelian; Sarma2015Majorana associated with the non-Abelian anyons enables nontrivial unitary operations among the degenerate many-quasiparticle states. Due to the fact that the qubits encoded in the global degenerate states are robust against the local systematic errors Witzel2006Quantum, the non-Abelian anyons can serve as the primitive units of topological quantum computation.
@@
Majorana zero mode (MZM), first theoretically predicted in the Kitaev chain model Kitaev2001Unpaired, represents the most elementary realization of the non-Abelian anyons. The Kitaev chain models a one-dimensional system of spinless fermions and was once regarded as a somewhat unphysical toy model Elliott2015Colloquium. Nevertheless, the spinless fermions can be simulated through experimentally accessible proxy systems, such as the spin-orbit-coupling electron gases under Zeeman field, as suggested by the subsequent theoretical studies Elliott2015Colloquium; Fu2008Superconducting; Sau2010Generic; Alicea2010Majorana; Lutchyn2010Majorana; Oreg2010Helical; Nadj2013Proposal; Pientka2013Topological; Gao2024Majorana. Experimental research for MZMs has been conducted in multiple physical platforms, including the superconducting systems Stevan2014Observation; Howon2018Toward; Jiao2018Chiral; Schneider2020Controlling; Jie2021Spin and hybrid magnetic-superconductor systems Schneider2021Topological; Schneider2022Precursors; Soldini2023Twodimensional. Recently, indirect evidences for MZMs have been witnessed in the hybrid semiconductor-superconductor system Microsoft2025Interferometric, exhibiting a zero-bias conductance peak Das2012Zerobias; Deng2012Anomalous; Finck2013Anomalous; Churchill2013Superconductor and the topological protection against the variations of local parameters Jiao2018Chiral; Jie2021Spin; Microsoft2025Interferometric. The braiding of MZMs, however, remains unimplemented as the gold-standard signature for definitively establishing their non-Abelian statistics Dvir2023Realization.
@@
Conventional braiding operations of MZMs were proposed under the adiabatic condition Ivanov2001NonAbelian; Stone2006Fusion; Read2009NonAbelian; Sau2011Controlling; vanHeck2012Coulomb; Kraus2013Braiding and governed by the Berry-Wilczek-Zee holonomy Wilczek1984Appearance; Cheng2011Nonadiabatic. In practice, the implementation of topological quantum computation based on the adiabatic braiding operations is constrained by the extended operation time. First, the prolonged exposure to the environment can induce ruinable decoherence effect on the quantum systems Carmichael1999statistical. Second, while the exact adiabatic solution is established within an infinite period of time, the realistic operation with a finite clock speed unavoidably entails nonzero nonadiabatic errors Nayak2008NonAbelian. Such errors will give rise to the leakage from the subspace of degenerate ground states to the external Hilbert space. Third, the long operation time of topological quantum computation increases the possibility of the quasiparticle poisoning Flsensberg2010Tunneling; Catelani2011Quasiparticle; Leijnse2011Scheme; Rainis2012Majorana; Karzig2021Quasiparticle; Liu2024Quasiparticle. The quantum information encoded in the braiding process of MZMs will leak out of the topological subspace due to the excitation of the extra quasiparticle, which can break the parity conservation of the MZMs subspace. Although the measurement-based tools Knapp2016Nature have been used to mitigate the unwanted transition errors during adiabatic braid operations Bonderson2008Measurement; Bomantara2020Measurement, a high-fidelity, robust, and fast braiding operation remains desired for MZMs.
@@
Inspired by transitionless quantum driving Berry2009Transition, the shortcuts to non-Abelian braiding Karzig2015Shortcuts; Wang2025Shortcuts exploit the time-dependent and direct exchange interaction between MZMs as the counterdiabatic driving fields. But practical implementation of such interactions in semiconductor-superconductor systems is fundamentally difficult Amundsen2024Colloquium. In analog to nonadiabatic holonomic quantum computation Sjoqvist2012Nonadiabatic; Liu2019Plug, the nonadiabatic braiding operations Yu2025Nonadiabatic have been proposed for a two-MZM system mediated by the lattice defect (LD). However, the braiding operations under the parallel-transport condition Jing2017Non is sensitive to the systematic errors in the local driving fields on LD.
@@
We here construct the braiding operations between any pair of three MZMs coupled to a mediator LD under local driving fields by the universal quantum control (UQC) theory Jin2025Universal; Jin2025Entangling; Jin2025ErrCorr; Jin2025Rydberg; Jin2025UniversalNon. It is a theoretical framework that unifies the shortcuts to adiabaticity and holonomic quantum transformation and is beyond the parallel-transport condition. The degree of freedom of LD can be eliminated Jin2025Rydberg; James2007Effective when the driving fields are largely detuned; and then the indirect exchange interaction among MZMs can be established. Consequently, the braiding operations of arbitrary two MZMs can be performed via the universal passages of the system. Assisted by the fast-modulated global phases, our braiding operations become insensitive to the significant fluctuations in both transition frequency of LD and Rabi frequency of driving fields and to the quasiparticle poisoning. In addition, the clockwise and counterclockwise population transfers of three MZMs can be demonstrated by the time evolution along the passage that is superposed by LD excited state and MZMs. Our protocol is also directly extendable when the mediator LD coupled to more MZMs.
@@
The remainder of this paper is structured as follows. In Sec. II.1, we introduce a general model that is consisted of three Kitaev chains coupled to the same lattice defect and the full Hamiltonian for LD and MZMs. In Sec. II.2, we derive the effective Hamiltonian with the largely detuned driving fields on LD, based on which the braiding operations can be established via universal nonadiabatic passages. In Sec II.3, we employ a rapidly varying global phase to correct systematic errors in driving fields, which gives rise to fault-tolerant, fast, and high-fidelity braiding operations for MZMs. In Sec. II.4, we evaluate our error-correction mechanism under the quasiparticle poisoning and discuss a combination strategy with the measurement-based braiding protocol. Section III is devoted to designing the chiral population transfer of three MZMs in the presence of LD excitation. The whole work is concluded in Sec. IV. Appendix A presents the UQC theory in the parametric space with an ideal setting, followed by the error-correction extension in Appendix B.
@@ II.1 Model and Hamiltonian
Figure 1: (a) Sketch of an atomic system comprising a lattice defect and three one-dimensional NN-site Kitaev chains. LD can be coupled to MZMs $\gamma_{1,1}$, $\gamma_{1,2}$, and $\gamma_{1,3}$ through the local off-resonant driving fields. Inset: the four-level transition diagram for the effective subspace of LD and MZMs. (b) Effective two-level transition diagram for the braiding operation between two MZMs $\gamma_{1,j}$ and $\gamma_{1,k}$, $j\neq k\in{1,2,3}$.
Our braiding protocol is conducted on the atomic-chain system as shown in Fig. 1(a). The system comprises a two-level system (lattice defect) as a mediator, consisting of an excited state $|e\rangle$ and a ground state $|g\rangle$, and three one-dimensional Kitaev chains Kitaev2001Unpaired; Kitaev2003Fault of open boundary condition, each of which is composed of $N$ sites. The system Hamiltonian can be written as Kitaev2001Unpaired; Kitaev2003Fault ($\hbar\equiv 1$)
$H(t)$ $=\sum_{k=1}^{3}\Big[-\mu\sum_{n=1}^{N}c_{n,k}^{\dagger}c_{n,k}-\sum_{n=1}^{N-1}(Jc_{n+1,k}^{\dagger}c_{n,k}$ (1)
@@
where $\sigma_{+}\equiv|e\rangle\langle g|$ ($\sigma_{-}=|g\rangle\langle e|$) and $c_{n,k}^{\dagger}$ ($c_{n,k}$) are the creation (annihilation) operators for LD and the $n$th site of the $k$th Kitaev chain, respectively. The fermions in the Kitaev chains are assumed to have the same transition frequency $\mu$. $\omega(t)$ is the transition frequency of the two-level lattice defect, which is set as a constant in our protocol despite it can be tuned via an external magnetic field in solid-state spin system Wolfowicz2021Quantum. $J$ and $g$ are the hopping and pairing coupling strength between neighboring fermions in the chains, respectively.
The Hamiltonian (1) admits an alternative representation in terms of Majorana fermion quasiparticles, which emerge as bounded pairs in the system. Specifically, the Majorana fermions of the $k$th Kitaev chain, $1\leq k\leq 3$, can be defined as Kitaev2001Unpaired; Kitaev2003Fault
@@
where $n$ runs from $1$ to $N$. As defined in Eq. (2), Majorana fermions are characterized by two fundamental properties: self-conjugation, $\gamma_{n,k}^{\dagger}=\gamma_{n,k}$ and anticommutation relations $\{\gamma_{n,j},\gamma_{m,k}\}=2\delta_{nm}\delta_{jk}$, where $n$ and $m$ run from $1$ to $2N$.
In the current hybrid semiconductor-superconductor systems Sau2012Realizing; Dvir2023Realization; Shi2024Probing; Bordin2025Enhanced; Nitsch2025Poor; Microsoft2025Interferometric, both LD and the fermions in the Kitaev chains are constituted by the quantum dots (QDs). The hopping coupling $J$ and pairing coupling $g$ between the neighboring quantum dots can be attained by elastic-to-tunneling and crossed Andreev reflection, respectively, mediated by a superconducting nanowire which hosts an Andreev bound state Liu2022Tunable; Bordin2023Tunable. Experimentally Sau2012Realizing; Dvir2023Realization; Shi2024Probing; Bordin2025Enhanced; Nitsch2025Poor; Microsoft2025Interferometric, both the coupling strengths $J$ and $g$ and the transition frequencies of LD $\omega$ and QDs $\mu$, can be individually manipulated by the gate voltages applied to the superconducting segments. $\omega/2\pi$ and $\mu/2\pi$ can be tuned within the range of $\sim(-10,10)$ GHz. Using the parametric setting of $\mu=0$ and J=gJ=-g Kitaev2001Unpaired; Kitaev2003Fault and Eq. (2), the Hamiltonian (1) is transformed to be
@@
It indicates that two decoupled Majorana zero modes appear at both ends of each Kitaev chain, i.e., $\gamma_{1,k}$ and $\gamma_{2N,k}$.
We desire to construct the nonadiabatic braiding operations between the chosen pair of MZMs $\gamma_{1,j}$ and $\gamma_{1,k}$, $j\neq k$, via a mediator LD. When the system is initially in the vacuum state $|0\rangle$, its dynamics is constrained in the subspace spanned by the states $|\gamma_{1,k}\rangle=\gamma_{1,k}|0\rangle$ in the absence of LD. In addition to the free Hamiltonian (3), we introduce the local time-dependent hopping interaction $J_{k}(t)$ and pairing interaction $g_{k}(t)$ between LD and the end fermion of three Kitaev chains Dvir2023Realization; Shi2024Probing; Bordin2025Enhanced; Microsoft2025Interferometric; Nitsch2025Poor. Then the full Hamiltonian can be expressed as Yu2025Nonadiabatic
@@
where the interaction or driving Hamiltonian reads,
$H_{d}(t)=\sum_{k=1}^{3}\left[J_{k}(t)\sigma_{+}c_{1,k}+g_{k}(t)\sigma_{+}c_{1,k}^{\dagger}+{\rm H.c.}\right].$ (5)
Similar to $J$ and $g$, the coupling strengths $J_{k}(t)$ and $g_{k}(t)$ can be effectively realized by the adiabatic elimination of the superconducting circuit in hybrid semiconductor-superconductor systems Dvir2023Realization; Shi2024Probing; Bordin2025Enhanced; Microsoft2025Interferometric; Nitsch2025Poor. In experiments, they are determined by the tunneling amplitudes for spin-conserving and spin-flipping processes and the BCS coherence factors and the eigenenergy of the Andreev bound state formulated in the superconductor circuit. Therefore, both $J_{k}(t)$ and $g_{k}(t)$ can become fully tunable via the gate voltages applied to the superconducting segments. Here it is required that $J_{k}(t)=g_{k}(t)=2\Omega_{k}(t)\cos[\int_{0}^{t}\omega_{k}(s)ds+\varphi_{k}]$. In typical experimental conditions, $J_{k}/2\pi$ and $g_{k}/2\pi$ can be tuned within the range of $\sim(-1,1)$ GHz Dvir2023Realization; Shi2024Probing; Bordin2025Enhanced; Microsoft2025Interferometric; Nitsch2025Poor. Consequently, the Rabi frequencies are about $\Omega_{k}(t)/2\pi\sim(-0.5,0.5)$ GHz. The Hamiltonian (5) has been extensively studied, both theoretically Liu2013Floquet; Li2014Tunable; Long2021Nonadiabatic; Nathan2021Quasiperiodic; Xu2023Dynamics and experimentally Cooper2019Topological; Sebastian2025Braiding, in cold atoms and quantum-dot arrays.
Adapting from MZMs, the full Hamiltonian $H_{\rm rot}(t)$ (4) suggests that there are no interactions between LD and the rest Majorana fermions $\gamma_{2n}$ and $\gamma_{2n+1}$, $1\leq n\leq N-1$. Then we can focus on
@@
$\times\cos\left[\int_{0}^{t}\omega_{k}(s)ds+\varphi_{k}\right]|e\rangle\langle\gamma_{1,k}|+{\rm H.c.}\Big\}.$
@@
where the detuning $\Delta_{k}(t)\equiv\omega_{k}(t)-\omega$. The relevant diagram of level transition can be found in the inset of Fig. 1(a). The two exponential terms in Eq. (7) have the same physical resource, i.e., the time-dependent driving frequencies $\omega_{k}(t)$. By tuning the magnitudes of $\Delta_{k}(t)$, the transition between the states $|\gamma_{1,j}\rangle$ and $|\gamma_{1,k}\rangle$ can be effectively engineered as the following subsection.